OpenStudy (anonymous):

Need help to solve this. ¼ (5x-1/2)-5/3<1/3

OpenStudy (anonymous):

thank you GT..i was confused on how to do that..

OpenStudy (mathteacher1729):

You can visualize this by going to http:www.wolframalpha.com and typing 1/4 (5x-1/2) < 1/3 + 5/3 The right side is just 6/3 = 2 So there are two equations: y = 1/4(5x-1/2) and y = 2 All values of x for which y = 1/4(5x-1/2) is LESS THAN (under) the line y =2 is a solution. :)

OpenStudy (anonymous):

thanks, will check out that site..

OpenStudy (anonymous):

Mathteacher, so the solution to that problem is NOT x>3/2 its x<17/10?

OpenStudy (anonymous):

Hey GT

OpenStudy (anonymous):

oh ok, just making sure, because with that graph it shows a different solution.

OpenStudy (mathteacher1729):

GT you made an error: 1/4(5x-1/2) -5/3 < 1/3 add 5/3 to both sides 1/4 (5x-1/2) < 1/3 + 5/3 realize that 1/3 + 5/3 = 6/3 = 2 AND distribute the 1/4 5x/4 - 1/8 < 2 ADD 1/8 to both sides... 5x/4 < 2 - 1/8 But you SUBTRACTED 1/8 from both sides... and everything below is in error. :( 5x/4 < 15/8 5x < 15/2 x < 3/2 (incorrect answer)

OpenStudy (anonymous):

mathteacher, i got x<17/10!!!???

OpenStudy (anonymous):

Actually, I made a mistake. I subtracted 1/8 instead of adding to the other side. Damn!!!

OpenStudy (mathteacher1729):

So the correct answer is x < 17/10 That is the correct answer. :)

OpenStudy (anonymous):

Sorry!

OpenStudy (mathteacher1729):

No prob gt. :) It happens.

OpenStudy (anonymous):

haha..thanks mathteacher..

OpenStudy (mathteacher1729):

Hint: when you are showing steps, always write a little next to each step explaining what it was that you did. :)

OpenStudy (anonymous):

1/4 (5x-1/2) < 1/3 + 5/3 5x/4 - 1/8 < 2 5x/4 < 2 + 1/8 5x/4 < 17/8 5x < 17/2 x < 17/10

OpenStudy (anonymous):

thank you both..